### v1

# Herding Unmasked: Insights into Cryptocurrencies, Stocks and US ETFs

Abstract
Herding behavior has become a familiar phenomenon to investors, carrying the potential danger of both undervaluing and overvaluing assets, while also threatening market stability. This study contributes to the literature on herding behavior by using a more recent dataset to cover the most impactful events of recent years. To our knowledge, this is the first study examining herding behavior across three different types of investment vehicle. Furthermore, this is also the first study observing herding at a community (subset) level. Specifically, we first explore this phenomenon in each separate type of investment vehicle, namely stocks, US ETFs and cryptocurrencies, using the widely recognized Cross Sectional Absolute Deviation (CSAD) model. We find similar herding patterns between stocks and US ETFs, while these traditional assets reveal a distinction from cryptocurrencies. Subsequently, the same experiment is implemented on a combination of all three investment vehicle types. For a deeper investigation, we adopt graph-based techniques such as Minimum Spanning Tree (MST) and Louvain community detection to partition the given combination into smaller subsets whose assets are most similar to each other, then seek to detect the herding behavior on each subset. We find that herding behavior exists at all times across all types of investment vehicle at a subset level, although the herding might not manifest at the superset level. Additionally, this herding behavior tends to stem from specific events that solely impact that subset of assets. Given these findings, investors can construct an appropriate investment strategy composed of their choice of investment vehicles they are interested in.
Keywords cryptocurrencies  $\cdot$ stocks  $\cdot$ US ETFs $\cdot$ herding behavior  $\cdot$ community detection
1 Introduction
Market participants have long been known to play a crucial role in shaping the price movements of financial assets [1]. High demand for an asset propels its prices upward, while a loss of interest or adverse news about it tends to lead to a price decrease. As a result, the understanding of investor reactions to various market conditions, e.g. financial crises, political events, bull and bear markets, is necessary to make wise investment decisions. Indeed, there have been a considerable number of studies focusing on herding behavior among investors in financial markets, from traditional ones like stocks and ETFs to alternative assets such as art real estate and crypto [2, 3, 4, 5, 6, 7, 8, 9]. Nowadays, this is still a topic of interest as financial markets are constantly changing and reacting to new events over time [2]. As defined by most studies, herding behavior describes collective actions from different investors, either purchasing or selling, targeting a set of different assets at a certain period which results in the same price movements between these assets [5]. This behavior typically arises from two scenarios [3]: First, intentional herding - where investors disregard their individual beliefs and follow the decisions made by the majority. Second, unintentional herding - where investors do not follow the market consensus and make a decision by themselves. However, in this case, they accidentally make the same decisions due to shared information sources such as financial reports, price momentum and analyses. It has been shown that the collective movements brought about by intentional and unintentional herding directly threaten market stability and are considered a key factor contributing to speculative bubbles [10].
Different approaches have been proposed to detect herding signals. These can be classified into statistical and Neural Network methods. The former often involves constructing a regression model based on return values, then using optimization techniques such as Ordinary Least Squares (OLS) to obtain independent features’ coefficients, which are assessed to identify the herding signals, using methods like Cross-Sectional Standard Deviation (CSSD model) [11], Cross-sectional Absolute Deviation (CSAD model) [12], Power Law Exponent-based model [4] and return-variance based model [5]. An alternative is to construct a metric based on a database of investors (e.g. the number of investors and demographics data) and their trading activities (e.g. volume of purchasing and selling), then set a threshold for the metric to decide whether the herding appears. This approach needs more than the return values since a database of such trading activities is required, such as measures proposed by Lakonishok et al. [13], Frey et al. [6] and Patterson et al.[14]. Regarding Neural Networks, these methods are rather new in this area. Currently, one mainly deploys Deep Learning models such as Long Short Term Memory (LSTM) [15] and BERT [16] to gauge the market sentiment (e.g. positive, negative and neutral) which is then incorporated as an independent feature in regression models akin to statistical approaches [17]. To this end, statistical approaches remain predominant in herding detection studies, with the CSAD model emerging as the most commonly utilized option [18, 19, 20, 21, 22].
One common finding agreed upon by most studies in different asset classes is that investors tend to follow the market consensus during periods of market uncertainty, which eventually results in herding behavior. Historical examples include events like Black Monday in 1987 [23], the Dot-Com/Tech crash in 2000 [23], the global financial crisis in 2007-2008 [24], and more recently the crash brought on by Covid-19 pandemic in 2020  [7]. Notably, it is worth mentioning that the role of algorithmic trading during market panics has been revealed, showing that this decision-making method still tends to operate during periods of high uncertainty [25]. However, this approach does not seem to reduce herding behavior. On the contrary, it appears to inadvertently cause herding due to identical decisions made by algorithmic trading systems [26]. In addition, studies in the cryptocurrency market indicate that its investors, who tend to be naïve, are more likely to behave irrationally and are easily influenced by price fluctuations as well as concrete items of news. By contrast, it has been suggested by a number of authors that professional investors tend to be dominant in traditional markets such as stocks, commodities and currencies. Besides, herding behavior in these markets is mainly explained by economic uncertainty as well as macroeconomic factors [8, 27, 4, 28].
Given the diversification of studies in herding behavior detection, most of them only focus on one particular type of investment vehicle. Since investors tend to diversify their portfolios by investing in different types of investment vehicle [29], only considering one of them overlooks this investment manner, resulting in a shortage of insights. Our study addresses this limitation by using three popular types of investment vehicle, e.g. stocks, US ETFs and cryptocurrencies.
Furthermore, existing literature tends to employ a herding detection model on the entire dataset comprising all available assets. As a result, conclusions regarding the herding phenomenon are obtained based solely on the whole dataset, potentially yielding inadequate results. For instance, the herding might just be true for a portion of the dataset while leaving the others unaffected but these latter are still counted because they belong to the same dataset. We propose an approach to tackle this limitation. In particular, we first use Louvain community detection [30] to group the assets based on their similarity in terms of price movements. Therefore, each group contains the assets that are most similar to each other on the basis of price movements. To recall, the fundamental idea of herding behavior is that assets display synchronized price movements due to the mutual trading actions of investors at the same time, especially during high market uncertainty. Thus, assets dissimilar to each other are less likely to demonstrate herding signals. To this end, our way of splitting assets can help reduce false positives in the results. Next, herding detection is applied to each group. Following this practice, we can observe the herding behavior at a community level, offering deeper insights compared to conducting herding detection on the entire dataset.
To address the shortcomings in the existing literature, our study aims to tackle the following research questions:
  • RQ1) Are herding patterns in each type of investment vehicle (namely stocks, US ETFs and cryptocurrencies) persistent, when considered separately? What causes the differences or similarities between the herding patterns?
  • RQ2) By forming a combination of all assets across three investment vehicle types and splitting them into small communities based on their similarity, are there patterns in herding that manifest themselves in this combination at a community level?
To answer these questions, we utilize 30-min closing price time series of 222 assets, including 146 stocks, 49 US ETFs tracking their corresponding US indices and 27 cryptocurrencies. All of the data is sourced from the platform FirstRate Data. As mentioned earlier, the CSAD model is one of the most popular approaches for herding detection at the moment, thus, we also use this model in our present work for ease of comparison with others’ work. Regarding community detection, we first construct an asset graph based on the Pearson correlation [31] between the assets, then apply Minimum Spanning Tree (MST) [32] and Louvain [30] algorithms to the asset graph in order to identify the communities. The details of this process are described in Section 3.
Results obtained from our study can be summarised into the following core findings: Firstly, while herding patterns in stocks and US ETFs are similar, there are differences between these and cryptocurrencies, meaning that these digital assets still remain separate from traditional assets to some extent. Additionally, the herding signal is stronger in the US ETFs compared to stocks and cryptocurrencies (RQ1). Secondly, by detecting herding behavior at a subset (community) level, we find that there are always subsets (communities) revealing herding behavior in each type of investment vehicle throughout the period examined, even though this phenomenon is not pronounced in the corresponding superset at some stages. Therefore, we conclude that while herding may not manifest when looking at all assets, it can still be found within specific subsets (RQ2). Thirdly, we notice 4 stock sectors (i.e. Information Technology, Finance, Healthcare and Energy), wherein assets within each sector and their corresponding ETFs tend to form communities themselves. Furthermore, the herding behavior observed in these communities tends to stem from specific events that solely impact that sector (RQ2).
The remainder of the article is organized as follows: Section 2 goes through related works. Section 3 provides a description of the datasets followed by methodologies. Section 4 describes the experimental results followed by implications and hypothesis. Lastly, the conclusion of this study is given in Section 5.
2 Related Works
2.1 Periods of Uncertainty in Recent Years
Throughout the 4-year span from 2019 to 2023, several significant events have hit global financial markets, including the trade war between China and the US [33], the outbreak of the Covid-19 pandemic [34], and, more recently, the political war between Russia and Ukraine [35]. While these events arose for different reasons, they collectively caused negative influences on financial markets, affecting both traditional and cryptocurrency assets to some degree.
The US-China trade war took place in January 2018, following the imposition of initial tariffs on Chinese imports by the new US administration at that time [33]. The primary goals behind this decision were to reduce the US trade deficits, stop the requirements of American companies to transfer technology and protect US technology companies from intellectual property theft [36]. As a countermeasure, China also levied different tariffs on the US products [33]. Although the trade war started early in 2018, it escalated during 2019 [37]. Consequently, financial markets in both the US and China were significantly affected, especially in the commodity, energy and technology sectors [38, 39, 37]. Major stocks in both countries as well as the ETFs that track them fell sharply [40, 41]. Trade war headlines attracted huge attention throughout the year. The US President Donald Trump’s announcement relating to the tariff imposition caused the stock and index values to fluctuate dramatically [42]. Indeed, it has been shown that there was a wave of selling pressure in 2019 because the stock prices fluctuated due to the panic of investors caused by the trade war [43]. In China, the economy also exhibited a decline in GDP growth rate in 2019 [36], accompanied by decreased returns across various sectors in financial markets [44]. The uncertainty in financial markets due to this trade war was also seen in other countries across the world [45, 39]. The event ended in January 2020 after both countries signed off on the so-called "Phase One agreement" trade deal [46].
Right after the trade war, the world encountered another challenge in the form of the Covid-19 pandemic [34]. This event disrupted economies across the world. In particular, due to the fast spread speed and serious consequence of this health crisis, different regulations were applied to stop the virus, such as closing national borders, constraining mobility as well as stopping business activities worldwide [47, 48]. Subsequently, these measures profoundly influenced the economies of numerous countries, eventually turning into a global economic downturn in March 2020 as a response to various government’s efforts to control the disease spreading [49]. This is considered one of the biggest financial crises ever, even worse than previous ones, e.g Black Monday 1987, 2007-2008 global financial crisis [50]. Various studies highlighted that this pandemic has disturbed and devalued financial markets, from traditional assets such as stocks and bonds [51, 52, 53] to safe-haven assets such as gold [54, 55] to the cryptocurrencies [51, 56, 57]. The pandemic’s strongest effects were observed during its first wave (from March until June 2020), with most assets experiencing a big drop in prices [58]. Fortunately, although the world continued facing different Covid-19 waves, its impact on financial markets lessened significantly, with their prices gradually recovering to pre-pandemic levels [59].
As the impacts of Covid-19 were being reduced through intervention measures such as vaccines and social distancing, another major event disrupted world economies. On 24 February 2022, Russia officially declared war on Ukraine. Numerous studies have dug into the impact of this event on the financial markets [60, 61, 35, 62, 63, 64]. Russia, a significant player in oil and gas production, has long been a crucial fuel supplier to many countries, particularly in the EU. Furthermore, Ukraine has served as a critical transit route for Russian petroleum exports to Europe. Consequently, this conflict had a severe effect on the energy sector [65]. In particular, when tensions rose and military actions escalated, oil prices tended to increase due to fears of potential supply disruptions. Conversely, when diplomatic efforts appeared to make progress or when there was a decrease in hostilities, oil prices tended to ease [66, 67]. Moreover, increased sanctions on Russia and efforts by countries, notably in Europe, to reduce dependence on Russian energy further triggered oil and gas shortages, the energy prices in Europe thereby surged during the first half of 2022 until alternative suppliers were found later on [68, 69]. Additionally, the EU’s embargo on seaborne crude oil and the G7’s price cap mechanism, which both took effect on December 5, 2022 also pressured the global oil prices already high for seasonal reasons [68]. Regarding other sectors, most studies have focused on the EU financial markets, since they were directly linked to Russia and Ukraine in commercial and political activities. Generally, it has been shown that the war had heterogeneous impacts on the EU financial markets. In particular, while most countries revealed a drop in their stock market prices after the war took place [62, 70, 63], those closest to Russia were impacted greater [71]. Interestingly, stocks and indices in several countries experienced an upward trend during the war, such as Belgium, Denmark, Luxembourg, Netherlands and Spain [62]. On the other hand, this war seemed not to have distorted the financial markets in Asian countries and other developed ones such as the US [64, 35, 72]. Nevertheless, studies [73, 72, 63] emphasized that, aside from the energy sector, the effect of this event on financial markets globally is negligible compared to previous events such as the Covid-19 pandemic, the trade war and past financial crises.
2.2 Herding Behavior in Traditional Markets
The discussion of herding behavior among investors dates back over a century, paralleling the long historical development of traditional markets [74]. One of the earliest topics discovered in this area deals with herding behavior under different market conditions. In [19], Aharon applied 2 popular herding detection methods, CSSD and CSAD, to a dataset comprising historical daily data spanning from 1990 to 2019. This dataset consists of 10 size-ranked portfolios constructed from stocks traded on the NYSE, AMEX and NASDAQ, covering different major events like the Asian crisis, the Dot-Com bubble, and the 2007-2008 global financial crisis. He found evidence of herding within these stock assets throughout the examined period. Moreover, by integrating the volatility index (VIX index) into the herding models, he noticed that herding signals intensified when market volatility - as measured by the VIX index - increased. The same conclusion was also reached by Ahn et al [4] who used different datasets and herding detection methods. Specifically, based on the daily US stock return time series of 137 firms listed in the Standard & Poor 500 (S&P500) from January 1992 to December 2021, they utilized the Power law exponent-based model to calculate the time-varying herding magnitude and then compared the results with an economic uncertainty index adopted from [75]. With access to longer and more recent data, the authors confirmed that, apart from the major events covered in the above study, the herding was also more pronounced during the Covid-19 pandemic.
However, one downside of these studies is that they detected the herding over an entire period, leading to a lack of details. Thankfully, several studies filled this gap by using shorter periods at different points in time, creating a more complete picture of this phenomenon. For example, Chiang and Zheng in [18] observed the herding of 63 stocks originally from Thailand between January 1997 and December 1998 - a period that covers the Asian financial crisis. Using the CSAD model and daily return time series, they suggested the presence of herding in these stocks during the crisis. On the other hand, in [76], the authors collected monthly prices from the 20 most traded stocks in the Egyptian market over 12 years from July 2002 until May 2014. As distinct from most studies, the authors used a state-space model to detect the herding while dividing the data sample into pre-crisis, during the 2007-2008 crisis and the post-crisis period. Their findings revealed an absence of herding behavior before the crisis, transitioning to herding during the financial crisis and persisting afterward.
Notably, with the huge effect of the Covid-19 pandemic, many studies have explored herding behavior during this unprecedented time [20, 77, 22, 78, 3, 79]. In [20], Bouri et al. selected 49 daily stock market indices representing 49 different countries sourced from Morgan Stanley Capital International (MSCI) indices. Using the well-known CSAD model and a rolling window with a size of 102 days, they examined the time-varying global herding from January 2019 until August 2020. Their idea was to assess whether financial markets around the world exhibited the herding phenomenon during the onset of the pandemic. Indeed, it was shown that herding was apparent from March to May 2020, coinciding with the worst period of the global economy caused by the pandemic. Other studies focusing on different types of investment vehicle, timeframes and methodologies also obtained the same results as investors tended to mimic each other’s investment decisions during the first wave of the pandemic (first and second quarter of 2020) [80, 2].
Another aspect explored in the study of herding behavior is the comparison between different countries. A work by [81] conducted a literature review about herding behavior in different countries from 1995 to 2021, its authors found differences in external factors driving herding within different regions. For instance, global risk factors were identified as primary culprits influencing the stock market and giving rise to herding behavior among investors in the Turkish market [82]. On the other hand, the Chinese market tended to be affected by factors such as analyst recommendations, short-term investors’ horizons as well as systemic risk [83]. Moreover, the authors also noticed that the majority of markets in developing countries were exposed to herding among retail investors (i.e. non-professional traders relying on personal judgment and subjectivity to make investment decisions). The purchasing power of these investors tends to be relatively small as limited by their personal earning ability. On the other hand, retail investors in developed countries tended to display less herding behavior due to the greater availability of information and knowledge of stock prices. Another study conducted by Chen and Zheng [84] employed the Simulated Method of Moments estimator to investigate the herding behavior in the Chinese and US stock markets. The experiment was based on the Shanghai Composite Index representing the China market and the S&P500 Index representing the US market over the period between 01/01/1992 and 10/30/2017, including 6739 daily observations. The authors revealed both similar and different patterns in herding behavior between the two countries. Furthermore, they stated that the Chinese stock market was mainly driven by behavioral sentiment dynamics due to the switching behaviors of investors while the US market was mainly driven by fundamental factors. The authors explained that this contradiction was attributed to the lack of a sufficiently transparent environment as well as comparatively few sophisticated investors in China, as compared to well-developed markets like the US. Additionally, over-speculative trading and the weak determining force of fundamental factors in China further contributed to this phenomenon. More recently, Yang and Chuang in [22] investigated herding behavior across three different countries, including China, Taiwan and the US. They collected stock returns for non-financial firms listed on the Taiwan Stock Exchange (TWSE), the Shenzhen Stock Exchange (SZSE), the Shanghai Stock Exchange (SHSE), and the New York Stock Exchange (NYSE) from January 2001 until June 2021. To observe the herding behavior, they combined the CSAD model with the Kalman filter and GARCH models. The authors discovered a relatively similar result to the above study. Specifically, Taiwan and China exhibited identical herding patterns with significant herding observed in two sub-periods: 2001-2004, coinciding with two local events, namely the Taiwan Strait crisis and the SARS pandemic and 2007-2009, corresponding to the global financial crisis. By contrast, although herding was also pronounced during 2007-2009 in the US, no herding was apparent before this sub-period. Instead, this developed market revealed another instance of herding during the Covid-19 pandemic in 2020.
In general, most studies in traditional markets found herding displays during events that disrupt financial and economic stability. As explained in [3, 79, 85], during such events, investors tend to abandon their personal information and are influenced by financial news as well as the actions of other investors. Moreover, investors in traditional markets tend to be easily influenced by word of mouth, given the increase in uncertainty and negative news surging, especially retail investors. Additionally, during uncertain times, both retail and institutional investors behave similarly and noise trading is shown to increase during the high volatility. On the other hand, institutional investors are dominant when the markets experience an upward trend, resulting in asymmetric herding during the positive market condition [4].
2.3 Herding Behavior in the Cryptocurrency Market
Cryptocurrencies are a relatively new asset class compared to traditional assets. Although the concept of digital coins was introduced in 2009 with the emergence of Bitcoin as the pioneering cryptocurrency [86], trading activities in this market only really commenced in 2013 [87]. As a result, studies on herding behavior in this domain mainly focus on the period from 2013 onward. In [88], Poyser identified herding in the top 100 cryptocurrencies by market capitalization from late April 2013 until mid-October 2019. He first used a Markov-switching model to segment the sample data into different regimes, then relied on a CSAD model to detect the time-varying herding in each regime. He found that herding behavior changed over time, with evidence of both herding and its absence throughout the examined period. Notably, herders were more active from mid-2017 until the end of the period since the signal of herding intensified during this interval. Additionally, in the second experiment, the author distinguished between 2 market conditions, namely positive and negative returns, to observe the herding behavior in each situation. The results revealed that cryptocurrency investors were more likely to follow the consensus when the market exhibited positive returns, while this tendency was diminished during periods of negative returns. Similarly, a study conducted by Raimundo et al. [89] collected more recent data, including 80 cryptocurrencies with the highest market capitalization from coinmarketcap.com between July 2015 and March 2020. In this work, they applied a state-space model adapting the standardized-beta method to detect the herding and used a regression model to test the relationship between the herding behavior and the market volatility. Like the previous study, the authors indicated a mix of herding and no herding throughout the considered period. However, they revealed a non-herding signal during most of 2019. Besides, they also found that the herding in cryptocurrencies became more pronounced when the market volatility increased. This finding tallies with various studies, using different datasets and methodologies [7, 90].
Similar to traditional markets, the Covid-19 pandemic has also attracted significant attention in the cryptocurrency market. In [7], the authors gathered intraday trading data for 9 large cryptocurrencies such as Bitcoin, Ethereum and Litecoin, covering a 2-year period from January 2019 to January 2021. They applied Patterson and Sharma’s herding intensity measure to each cryptocurrency in two distinct sub-periods: Pre-Covid-19 (January 2019 - December 2019) and Covid-19 pandemic (January 2020 - January 2021). Unlike other aforementioned approaches, this method can detect the herding activities in each asset. Results suggested a dramatically greater herding intensity during the latter sub-period across all examined cryptocurrencies. Nevertheless, while all cryptocurrencies showed the herding phenomenon during this time, no evidence of herding was found in Bitcoin, Ethereum, and Ripple during the Pre-Covid sub-period. These findings align with those of Youssef and Waked who executed a hybrid model combining CSAD and Markov-switching methods on 43 cryptocurrency price time series with daily frequency [21]. In particular, their dataset spans from 28/04/2013 to 11/11/2020, divided into two sub-samples based on the Covid-19 pandemic, including the Pre-Covid from the beginning until the end of 2019 and the Covid outbreak thereafter. As expected, only the second sample (i.e. the Covid outbreak) revealed the herding evidence. Furthermore, for the second part of the work, the authors used the coronavirus media coverage index (MCI) which calculates the percentage of all news sources covering the novel coronavirus topic to test the relationship between news and the herding magnitude. Interestingly, they found that the news about the pandemic significantly drove the herding behavior among investors in the cryptocurrency market, indicating the fact that cryptocurrency traders are mainly influenced by news and people’s sentiments. The result is in line with most of the existing literature [91, 92, 93, 94].
More recently, a few studies have utilized a newer dataset that covers up to the year 2023 [95, 96]. However, instead of segmenting the sample into smaller intervals like the previous studies did, they mainly retained the whole period which caused ambiguity for herding detection models. Consequently, their findings tended to suggest an absence of herding throughout the examined period, even during highly uncertain times such as the outbreak of the Covid-19 pandemic.
From the existing literature, it is evident that most studies investigating the financial markets (both traditional and cryptocurrency) focus on periods before the year 2022, while only a few works extend into more recent times. To this end, our work contributes to the existing literature by using more up-to-date data, which spans from April 2019 until May 2023. Moreover, although several studies divided their data into smaller sub-periods, these divisions seem not to be sufficiently refined. To address this, we segment our data into finer sub-periods corresponding to major events that occurred in the financial markets throughout the examined period. Further details of this time division will be described in Section 3.
3 Methodology and Data Collection
3.1 Data Sources
The historical time series utilized in this study are collected from the FirstRate Data, a leading provider offering high-resolution intraday prices of financial assets on a variety of asset classes. Our dataset comprises 222 assets categorized into 3 types of investment vehicle, namely cryptocurrency (27 assets), stock (146 assets) and the US ETF (49 assets). The selection of these assets is based on four criteria:
  1. i.
    Data availability: the assets available on the FirstRate platform.
  2. ii.
    Market capitalization: since we are limited to little more than 50 assets for cryptocurrencies and US ETFs, we try to use all of them. On the other hand, there are thousands of stocks available to use so we narrow these down to the top 200 stocks whose companies are in the top 200 in terms of market capitalization, according to the web page Companies Market Cap.
  3. iii.
    High trading frequency: To avoid missing values, we only use assets with high trading frequency. As a result, we have to remove the assets that have more than 1, 10 and 12 percent of missing values among stocks, cryptocurrencies and US ETFs, respectively. Indeed, after filtering the assets, the majority of the remaining have less than 1 percent of missing values.
  4. iv.
    In the case of US ETFs: the ability to act as a proxy for a corresponding US stock index.
Each time series covers the period from 01/04/2019 until 03/05/2023 and has a resolution of 30 minutes. Since the official trading time of traditional markets is between 09:30 a.m. and 4:00 p.m., we select this timeframe as the boundary for our time series, meaning that the cryptocurrency data is only available within this trading time, even though these coins are traded nonstop. Other price values outside of this timeframe are not considered in our experiments. This practice facilitates the comparison between assets from different types of investment vehicle and also preserves their equality. A list of assets in each type of investment vehicle is given in Appendix B.
3.2 Financial Sector Classification
Unlike cryptocurrencies and US ETFs, the stock market is classified into 11 sectors based on their respective business activities: Communication Services, Utilities, Real Estate, Materials, Information Technology, Industrials, Healthcare, Financials, Energy, Consumer Staples, and Consumer Discretionary. In addition, while there is no official classification for ETFs and most contain assets from different stock sectors, we notice the existence of ETFs designed to track the performance of a specific stock sector. For instance, ticker IYW represents the iShares U.S. Technology ETF which tracks technology stocks. Thus, although we do not classify ETFs, we keep in mind this characteristic. As a result, we assign each asset to a specific sector, either cryptocurrency, US ETF, or one of the aforementioned stock sectors. We use this information in our experiments for a deeper investigation and analysis of the grouping of such assets.
We note that the number of assets in each stock sector is different and depends on its asset’s market capitalization. For instance, Technology, Finance and Healthcare stocks tend to have very high market capitalization while Communication Services and Real Estate tend to have lower market capitalization [97]. Consequently, among the 11 stock sectors, Utilities, Materials, Real Estate and Communication Services exhibited a limited number of entities. Therefore, this study concentrates only on the remaining 7 sectors. Further details for this classification are provided in Appendix C.
3.3 Graph Construction and Minimum Spanning Tree
A graph is a set of objects (referred to as nodes) that show connections with each other. That is, if there is a connection between two objects, it is marked by a link (referred to as an edge) connecting them together. The definition of this connection depends on the purpose of the graph. For instance,  [98] constructed an electric bicycle-based commute graph whose nodes represent locations and the connection between two nodes is measured by the number of trips traveling between the two corresponding locations that use this kind of bicycle. On the other hand, the Belgian mobile phone graph was introduced in [99] where the authors examined the connections among phone users via their phone calls and text messages’ frequency. Regarding financial markets, a graph between financial assets tends to focus on the correlation between them, i.e. the similarity in price movements between the assets [100, 56, 58, 57]. In this study, our graph is constructed as follows.
Given a price time series $x_{i}=(x^{1}_{i},x^{2}_{i},\dots,x^{T}_{i})$ where $i$ represents an asset, $T$ represents the length of the time series, $x^{t}_{i}$ represents the price value of asset $i$ at time $t$. The log-return at time $t$ of asset $i$ is defined as $r^{t}_{i}=log(r^{t}_{i}/r^{t-1}_{i})$. Although the price values are undoubtedly informative, we use the return value for this study as it is represented as a percentage, making them scale-free and especially, stationary, which is an important requirement for statistical tools, such as herding detection models [101].
To obtain a graph of assets, we first construct a correlation matrix $\mathbf{C}$ based on the similarity between the assets. In particular, each element $c_{ij}$ in $\mathbf{C}$ represents the similarity between 2 assets $x_{i}$ and $x_{j}$ such that
$c_{ij}=\frac{\sum\limits_{t=1}^{T}(x_{i}^{t}-\bar{x_{i}})(x_{j}^{t}-\bar{x_{j}})}{\sqrt{\sum\limits_{t=1}^{T}(x_{i}^{t}-\bar{x_{i}})^{2}\sum\limits_{t=1}^{T}(x_{j}^{t}-\bar{x_{j}})^{2}}}$ (1)
Where $\bar{x_{i}},\bar{x_{j}}$ are the average values of time series $x_{i},x_{j}$, respectively. This similarity measure is known as the Pearson correlation measure [31] and ranges from -1 to 1, where -1 indicates a total negative linear relationship, i.e. two time series are completely different, 0 indicates no linear correlation, and 1 indicates a total linear relationship, i.e. two time series are completely similar.
We then determine a distance matrix $\mathbf{D}$ where each element $d_{ij}=\sqrt{2(1-c_{ij})}$. This distance matrix transformation is necessary as it satisfies the characteristics of a metric [102] that are necessary for further steps. Essentially, $d_{ij}$ represents the distance between two time series $x_{i}$ and $x_{j}$, ranging from 0 to 2 in which 0 indicates a complete similarity and 2 indicates a complete difference between them.
From the distance matrix, we construct an asset graph where each node is an asset and an edge between two nodes depicts the relationship between them, which is the similarity calculated in the previous steps. One downside of the graph is that it is dense, making it memory-intensive and time-consuming for processing because all nodes are linked together while there are edges with negligible information. As a result, it costs a considerable amount of time to process the graph and redundant information might confuse experimental results. Thus, we extract a subset of the graph called Minimum Spanning Tree (MST)[32] to reduce the graph size and remove the redundant information. This subset keeps edges with the highest correlation degree, guarantees the characteristics of a tree and ensures other edges are removed. Hence, only the important part of a graph is kept. As a result, this technique reduces our graph size of 222 nodes and 24531 edges to 222 nodes and 221 edges.
3.4 Louvain Community Detection
Community detection aims to partition nodes of a graph into separate groups (called communities) by maximizing the similarity between nodes within a community while maximizing the distinction between different communities. In other words, this detection results in communities where nodes belonging to a community are most similar to each other while nodes from different communities exhibit different characteristics. For this technique, Louvain community detection algorithm is well-known for its applications in large networks [30]. Additionally, this approach is rather straightforward and is also adopted widely in various research topics [103, 104, 105]. Since our graph has more than 200 nodes, this method is unquestionably well-suited for this case.
The Louvain algorithm is based on modularity optimization and its fundamental principle is searching for a community partition such that it maximizes the density of edges inside communities in comparison with edges between different communities. This optimization is executed through maximizing a measure called modularity:
$Q=\frac{1}{2m}\sum\limits_{i,j}\left[A_{ij}-\frac{k_{i}k_{j}}{2m}\right]\gamma\left(g_{i},g_{j}\right)$ (2)
Where $A$ represents a graph, $A_{ij}=1$ if there is an edge between node $i$ and node $j$ and 0 otherwise, $m=\frac{1}{2}\sum\limits_{i,j}A_{ij}$ is the total number of edges in the graph, $k_{i}=\sum\limits_{j}A_{ij}$ is the node degree of node $i$, $g_{i}$ represents the community where node $i$ belongs to, $\gamma\left(g_{i},g_{j}\right)=1$ if $g_{i}=g_{j}$ (nodes $i$ and $j$ belong to the same community) and 0 otherwise. To this end, the best community partition is obtained when $Q$ is maximum.
From this idea, the process of Louvain community detection comprises two phases, as shown in Pseudocode 1. Moreover, Fig 1 illustrates the process of constructing communities from a set of time series.
Input: Graph $A=(V,E)$ with $V$ as the set of nodes and $E$ as the set of edges.
Output: Set of communities
Initialize: Each node in the graph is assigned to a separate community.
repeat
       Phase 1;
       while modularity $Q$ can be improved do
             for each node $i\in V$ do
                   Calculate the modularity gain $\Delta Q$ in turn for moving $i$ from its current community to each neighboring community;
$\Delta Q=\left[\frac{\sum_{in}+2k_{i,in}}{2m}-\left(\frac{\sum_{out}+k_{i}}{2m}\right)^{2}\right]-\left[\frac{\sum_{in}}{2m}-\left(\frac{\sum_{out}}{2m}\right)^{2}-\left(\frac{k_{i}}{2m}\right)^{2}\right]$
Where $\sum_{in}$ is the number of edges between nodes inside the community $g_{j}$, $\sum_{out}$ is the number of edges linked to nodes inside $g_{j}$ (including edges from nodes within other communities), $k_{i,in}$ is the number of edges from node $i$ to nodes in $g_{j}$;
                   if $\Delta Q$ in all cases is not positive then
                         $i$ stays in its current community;
                  else
                         Move $i$ to the community where $\Delta Q$ is positive and maximum;
                   end if
             end for
       end while
      if modularity $Q$ after Phase 1 increases then
             Phase 2;
             Construct a new graph $A^{\prime}$ where each community found in Phase 1 is considered as a node of $A^{\prime}$;
             The weight of an edge between two new nodes is the number of edges between nodes in the two corresponding communities;
             $A\leftarrow A^{\prime}$;
      else
             Stop the algorithm;
       end if
until No further improvement in modularity;
Algorithm 1 Louvain Community Detection
Figure 1: The process of implementing community detection from a set of time series. For more information, see Clarification 1..
3.5 Herding Detection Strategies
Cross-Sectional Absolute Deviation (CSAD) model is one of the most well-known herding behavior detection approaches that has been used widely in many studies[90, 106, 95, 20, 85]. The fundamental idea of this technique is to examine the relationship between the average return (market return) of a set of assets and the dispersion across each asset return (CSAD). A herding signal is pronounced when the market return changes significantly (i.e. the market experiences large fluctuations, either going up or down) while the dispersion between asset returns decreases (i.e. the return movements of different assets become more similar). In this study, we use two forms of the CSAD model for our experiments.
$CSAD_{t}=\beta_{0}+\beta_{1}|r_{m,t}|+\beta_{2}r_{m,t}^{2}+\epsilon_{t}$ (3)
$CSAD_{t}=\gamma_{0}+\gamma_{1}D^{up}|r_{m,t}|+\gamma_{2}D^{up}r_{m,t}^{2}+\gamma_{3}D^{down}|r_{m,t}|+\gamma_{4}D^{down}r_{m,t}^{2}+\zeta_{t}$ (4)
Where $CSAD_{t}=\sum\limits_{k=1}^{T}|r_{i,t}-r_{m,t}|/T$; $r_{i,t}$ is the return of asset $i$ at time $t$; $r_{m,t}=\sum\limits_{k=1}^{N}r_{k,t}/N$ is the average of asset returns (also known as the market return) at time $t$; $\beta_{0},\beta_{1},\beta_{2},\gamma_{0},\gamma_{1},\gamma_{2},\gamma_{3},\gamma_{4}$ are regression coefficients, $D^{up}=1$ if $r_{m,t}$ is positive, otherwise 00; $D^{down}=1$ if $r_{m,t}$ is negative, otherwise 00; $\epsilon_{t},\zeta_{t}$ are error terms at time $t$.
Regarding Eq 3, a herding exists when the parameter $\beta_{2}$ is significantly negative (a negative coefficient with a confidence level of at least 90%). Intuitively, we expect this coefficient to be negative because it follows the idea of the herding behavior, i.e. a dramatic increase or decrease in the market return ($r_{m,t}^{2}$) coincides with a decline (negative $\beta_{2}$) in the dispersion between corresponding assets’ returns ($CSAD_{t}$). Eq 4 refers to a more detailed version of Eq 3 in which we separate the market condition into up (i.e. positive market return) and down (i.e. non-positive market return) so that the herding can be detected during the up and down market condition separately, instead of considering all market conditions (i.e. both up and down equally) as declared in the first form. The use of both forms for herding detection comes from an expectation that there might be periods when herding does not exist when considering both up and down market conditions at the same time but it might be found in the up or down market condition only. To this end, the same logic of Eq 3 is applied to Eq 4 wherein a significantly negative $\gamma_{2}$ refers to a herding signal during the up market condition while a significantly negative $\gamma_{3}$ refers to a herding signal during the down market condition.
3.6 An Early Experiment: Herding Behavior throughout the Entire Period
We start this study by performing all experiments on the whole examined period without portioning it out. Therefore, based on our research questions, herding behavior detection for the entire period is applied to each type of investment vehicle(i.e. cryptocurrencies, stocks and US ETFs) followed by a detection on the combination of these three types of investment vehicle at a community level.
Table 1 provides the estimation of $\beta_{2}$ from Eq 3 and $\gamma_{2},\gamma_{3}$ from Eq 4 for each type of investment vehicle. All of them reveal signs of herding behavior over the considered timeline from April 2019 until May 2023 as the coefficient $\beta_{2}$ is statistically negative in all cases, even though their herding magnitudes are different. When distinguishing the market condition into up and down, herding is pronounced in both conditions for stocks and US ETFs, however, only the down condition in the cryptocurrency market exhibits this phenomenon.
[…]
Coefficient Crypto US ETF Stock
$\beta_{2}$1.751* -0.859* -3.023*
$\gamma_{2}$
(up market)
3.465* -3.100* -1.354*
$\gamma_{3}$
(down market)
-2.258* -2.766* -0.854*
The parameter $\beta_{2}$ is estimated from the equation:
$CSAD_{t}=\alpha+\beta_{1}|r_{m,t}|+\beta_{2}r_{m,t}^{2}+\epsilon_{t}$.
The parameters $\gamma_{2}$ and $\gamma_{3}$ are estimated from the equation:
$CSAD_{t}=\alpha+\gamma_{1}D^{up}|r_{m,t}|+\gamma_{2}D^{up}r_{m,t}^{2}+\gamma_{3}D^{down}|r_{m,t}|+\gamma_{4}D^{down}r_{m,t}^{2}+\zeta_{t}$.
Fig 2 illustrates herding detection results and sector distribution of each community (as a result of applying the Louvain community detection algorithm) in the combination of all three types of assets over the entire period. Specifically, the number of communities found by the Louvain algorithm and the proportion of each sector in each community is illustrated in the lower sub-figure. Sectors are listed on the y-axis, each represented by a unique color. For each community, the longer the bar of a corresponding sector, the greater the sector’s percentage within that community. In the meantime, the upper sub-figure reveals the corresponding herding detection results for each community in the lower sub-figure. The grey bar means the herding based on CSAD model (Eq 3) is significant, the green bar means the herding based on CSAD for the up market condition (Eq 4) is significant, the red bar means the herding based on CSAD for down market condition (Eq 4) is significant. The higher the bar, the more pronounced the herding. The assets in each community are listed in Appendix D. We find that most communities reveal herding except for the two communities 7 and 8. Overall, it is shown that performing herding detection on the whole period tends to finish with a ’Yes’ result. However, it is almost impossible to conclude anything given this result since the examined period is long, leading to a lack of detail. Furthermore, all types of investment vehicle exhibit herding behavior so no unique patterns can be found. This ambiguity encourages us to split the original period into sub-periods so that the herding behavior can be examined on a finer time interval, which enhances the detail level of the experiment. The information for this timeline division is proposed in the next section.
(a) Herding Detection Results
(b) Sector Distribution
Figure 2: Herding detection on each community (a) and its sector distribution (b) over the entire period from 01/04/2019 until 03/05/2023. More information on each sub-figure can be seen at Clarification 2..
3.7 Timeline Division
As mentioned earlier, our dataset covers a 4-year period from 01/04/2019 to 03/05/2023. Throughout these years, the financial markets experienced several pressing events, including the US-China trade war [38, 39, 45], the Covid-19 pandemic [56, 53, 50], and the Ukraine-Russia conflict [60, 61, 35, 62]. Since different events are characterized by different behavioral trends/reactions of investors, we split the given period into smaller intervals based on the events that occurred. Following this, herding detection-related experiments in this study are run on each sub-period. This practice helps us track the changes in herding behavior over time as well as recognize common features appearing during a particular event.
Notably, after Covid-19’s first wave, the financial markets began to enter the recovery time and most assets in both traditional and cryptocurrency markets have witnessed a surge in prices since the beginning of 2021. This was a period when most assets reached their all-time highest values. Since 2022, the markets have calmed down gradually, marking the end of the bullish time in April 2022 for most assets. Thus, we also include this Bull market time in our time division. Table 2 displays an explicit division for the original period.
Table 2: Five sub-periods declaration.
Name Pre-Covid-19 Covid-19 pandemic Post-Covid-19 Bull market
Ukraine-Russia
conflict
Time interval
01/04/2019
to
31/12/2019
01/01/2020
to
30/06/2020
01/07/2020
to
31/12/2020
01/01/2021
to
30/04/2022
01/05/2022
to
03/05/2023
4 Experimental Results and Discussion
This section investigates the herding behavior in cryptocurrencies, US ETFs and stocks, across different sub-periods. The results are then synthesized to explore insights about this phenomenon. Our experiments are structured into two main tasks according to the defined research questions: Firstly, we implement herding detection in each separate type of investment vehicle (RQ1). Secondly, we combine all three aforementioned types of investment vehicle and perform herding detection on the combination (RQ2). This idea comes from the hypothesis that the view on each type of investment vehicle differs from a composition of multiple different ones. In particular, knowing the herding patterns of a specific market is important for understanding its unique features, major events that the market has experienced, and which kind of information/news impacts the market, etc. On the other hand, observing different types of investment vehicle at the same time is critical for understanding the connections between them, which can highlight the way investors diversify their portfolios or the impact of each separate type of investment vehicle on their combination. Such insights cannot be obtained by focusing only on a single market.
4.1 Herding Detection on Each Type of Investment Vehicle
In this part, we tackle the first research question by uncovering herding signals in each type of investment vehicle across the five aforementioned sub-periods. Subsequently, answering various sub-questions such as: Which events cause the phenomenon? How strong is it? How are investors’ reactions linked to the herding? How resilient is the herding? Then, a comparison between the three types of investment vehicle is conducted.
Table 3 provides the estimation of $\beta_{2}$ from the CSAD model (without distinguishing up and down market conditions) while Table 4 provides the estimation of $\gamma_{2}$, and $\gamma_{3}$ from CSAD for the up market condition and CSAD for the down market condition, respectively, for each type of investment vehicle. Only values with the star symbol are statistically significant. Herd behavior is revealed at a period only when its corresponding coefficient (either $\beta_{2}$, $\gamma_{2}$ or $\gamma_{3}$) is statistically significantly negative. The more negative the value, the more pronounced the herding.
[…]
CSAD ($\beta_{2}$)
Periods Crypto US ETF Stock
Pre-Covid-19 0.0801* -4.550* -0.0894*
Covid-19 pandemic -0.0416* -0.995* -0.0281*
Post-Covid-19 -0.0524* 8.684* 0.1032*
Bull market -0.0271* 31.398* 0.0265
U-R conflict 0.0561* 0.155 0.0144
The parameter $\beta_{2}$ is estimated from the equation $CSAD_{t}=\alpha+\beta_{1}|r_{m,t}|+\beta_{2}r_{m,t}^{2}+\epsilon_{t}$.
[…]
Up Market Condition ($\gamma_{2}$) Down Market Condition ($\gamma_{3}$)
Periods Crypto US ETF Stock Crypto US ETF Stock
Pre-Covid-19 0.1133* -2.064 0.0001 0.0585* -5.085* -0.0951*
Covid-19 pandemic -0.0379* -1.297* -0.0305* -$0$998* -0.0270*
Post-Covid-19 -0.0471* 9.528* 0.0774* -0.0474* -0.330 0.0988*
Bull market 0.0229* 6.725* 0.0603** -0.0276* 33.658* 0.0262
U-R conflict 0.0511* 0.464 0.0198 0.0604* -0.417 0.0049
The parameters $\gamma_{2}$ and $\gamma_{3}$ are estimated from the equation
$CSAD_{t}=\alpha+\gamma_{1}D^{up}|r_{m,t}|+\gamma_{2}D^{up}r_{m,t}^{2}+\gamma_{3}D^{down}|r_{m,t}|+\gamma_{4}D^{down}r_{m,t}^{2}+\zeta_{t}$.
Starting with the cryptocurrency market, the herding coefficient $\beta_{2}$ reveals herding signals in three sub-periods, including during the Covid-19 pandemic, the Post-Covid-19 and the Bull market time. Surprisingly, the strongest herding is found in the Post-Covid-19 while the peak of the Covid-19 pandemic, which we expected to show the strongest herding behavior, only comes in second place. On the other hand, the Pre-Covid-19 and Ukraine-Russia conflict sub-periods disclose an absence of herding behavior. The results remain mostly unchanged after separating the market into up and down conditions. Specifically, the coefficients $\gamma_{2}$ and $\gamma_{3}$ are only statistically significantly negative during the sub-periods when $\beta_{2}$ is negative, indicating that either the up or down market condition or both conditions reveal the evidence of herding in these stages. However, while herding is pronounced in both up and down market conditions during the Covid-19 pandemic and Post-Covid-19 sub-periods, only the downturn times in the Bull market time have the appearance of herding. These results sufficiently capture the prevailing sentiment among cryptocurrency investors. Specifically, they seem to be prone to be influenced by market fluctuations and breaking news on the internet [107], leading to similar investment decisions, i.e. selling off their investments when the market experiences a downturn or bad news circulates on the internet to avoid a loss while purchasing new investments during bullish phases or when positive news about cryptocurrencies spreads among the general public due to the Fear of Missing Out (FOMO) on an increase in the price of the coins[90]. The high degree of sensitivity of investors in this digital market may be attributed to the dominance of irrational (naïve) participants, as confirmed in different studies [8], these people participate in the cryptocurrency market primarily because of its great potential for huge profit despite their lack of investment experience. Therefore, they tend to ignore their own analysis and information by following the decisions made by the public due to their lack of financial knowledge [27]. A greater herding magnitude during the Post-Covid-19 compared to the peak of the pandemic can be attributed to the cryptocurrency market’s tendency to be more influenced by positive circumstances than negative ones [108, 19]. During this time, the market started to recover from the decline caused by the pandemic, experiencing a price surge since the fourth quarter of 2020. An upward trend in prices along with positive word of mouth, as well as financial news, triggered the positive sentiment of investors together with the FOMO phenomenon that eventually caused a purchase herding in the market [27, 96, 21]. When it comes to the Bull market time, although the herding is evident, it appears to be weaker than other sub-periods. This might be because many investors had already purchased the coins at the earlier stage and were more likely to hold onto those coins during this time because of the positive expectations of an upward trend in the future. Thus, they barely purchased at this time, instead, the herding might come from them executing a sell action. Indeed, we notice that there was a downturn phase in this sub-period when the market significantly shrank. Hence, it is possible that a portion of investors began selling off their investments as they either felt panicked by the price drop or were motivated to take profits, resulting in a period of herding [90, 106, 109, 90]. Our result when examining the cryptocurrency market in up and down conditions also shows that only $\gamma_{3}$ is negative during this phase, meaning that only the downturn time in the Bull market time reveals the herding behavior, which reinforces our hypothesis. Overall, our results of herding detection are in line with existing literature which also demonstrates evidence of herding during the first outbreak of the pandemic/ 2020 global economic crisis and the periods of bullish markets, while there was no herding signal found before the pandemic and after the bullish time. [7, 21, 96, 110].
The herding behavior observed in the stock and US ETF markets displays similar patterns, with only the first two sub-periods, namely Pre-Covid-19 and Covid-19 pandemic, having signs of herding behavior. These time intervals cover critical events that significantly impacted traditional markets, i.e. the US-China trade war during the year 2019 and the severe hit of the contagious health disease Covid-19 on the global economy and financial markets in the first half of 2020. Notably, the US-China trade war took place throughout 2019 and continually affected the economies of not only the US and China but also other countries and continents [45, 39]. Moreover, investors have also been bombarded by both negative and positive news about this event, with the largest contribution coming from President Trump who often used social networks to post updates on the trade war. Consequently, such news dragged the markets up and down frequently [42]. Indeed, it has been shown that there was a selling wave of assets during this year [43]. This explains the more pronounced herding within the negative market return compared to the positive market return during the Pre-Covid-19, in both stock and US ETF markets.
The herding in the stock and US ETF markets delivers two interesting results. Firstly, herding is more pronounced during the Pre-Covid-19 compared to the Covid-19 pandemic sub-period. Therefore, although the pandemic led to greater losses and fluctuations in traditional financial markets, it is noteworthy that investors exhibited more intense herding behavior before the pandemic, particularly during the turbulent time caused by the US-China trade war. One possible reason for this is the constant fluctuations of markets throughout 2019 when the trade war became more and more intense. In particular, different tariffs and quotas were imposed between these two countries throughout this year, directly impacting financial markets worldwide. As a result, the constant release of negative and positive news gave rise to frequent market movements. Thus, investors actively reacted to market situations, causing herding behavior on a regular basis. The impact of the Covid-19 pandemic was truly severe and damaging within the second quarter of 2020 but this was followed by a gradual recovery. Hence, investors panicked during the peak of the pandemic but things quickly got back to normal, and as they started to make objective investment decisions again, herding then lessened. Another possible reason is that investors in traditional markets are more influenced by macroeconomic factors as well as policy uncertainty. Specifically, US interest rates increased because the government expected the economy to grow in 2019. Furthermore, the unemployment rate increased due to the trade war and GDP in 2019 also fell to its lowest since the 2008 crisis [111, 112, 4]. These macroeconomic factors had the potential to scare traditional market traders.
Secondly, the magnitude of herding behavior in the market composed of US ETF appears to dramatically surpass that observed in the stock market (and also the cryptocurrency market). We explain this distinction by examining the CAPM beta score [113] of each asset. An asset’s beta score measures its returns’ tendency to follow the overall market. In other words, it measures how likely an asset’s returns increase (or decrease) when the overall market’s returns increase (or decrease) where the overall market tends to be the S&\&&P 500 index. A beta of unity means that the asset’s movements are exactly the same as the overall market while less than unity indicates weaker movements, resulting in a lower return compared to the overall market and greater than unity indicates stronger movements, resulting in a higher return compared to the overall market. From Fig 3, we find that beta scores of US ETFs tend to be close to each other and show a tight range over those of the overall market. On the other hand, the beta scores of stocks show a wider dispersion. However, all of these traditional assets exhibit a correlation with the overall market to some extent as their beta scores are close to unity. By contrast, cryptocurrencies exhibit a different behavior since not all of them follow the movements of the S&\&&P 500 index, indicated by the beta scores of almost zero. Instead, cryptocurrencies tend to follow the movements of Bitcoin. Additionally, to compare the beta scores between stocks, US ETFs and cryptocurrencies numerically, we also calculate the average distance of assets’ beta scores in each type of investment vehicle compared to the overall market as measured by the S&\&&P500. We find that the beta scores of US ETFs are closer to unity, as compared to stocks and cryptocurrencies which reveal a higher distance value compared to US ETFs. This result is shown in Table 5. Therefore, US ETFs have a higher tendency to follow the movements of the overall market (S&\&&P500) when compared to underlying stocks and cryptocurrencies, which explains the high magnitude of the herding behavior of US ETFs. Another possible reason might be the high liquidity of these stock ETFs. Specifically, because the liquidity of ETFs is higher than stocks [114, 115], their prices mirror investors’ reactions (e.g. buy or sell) quicker [116]. Consequently, the chance for their price movements to become alike is higher than that of underlying stocks, resulting in stronger herding. On the other hand, stocks tend to have lower liquidity [114, 115] and the degree of liquidity across different stocks might be very different. Thus, even if investors react similarly to different stocks at the same time, the difference in different stocks’ liquidity causes a lag between their price movements and lessens the herding.
Figure 3: Box Plots of Beta Scores in each Market. The overall market is the S&\&&P 500 index for traditional assets (i.e. stocks and ETFs) while both S&\&&P 500 index and Bitcoin are used as the overall market for cryptocurrencies.
Table 5: Average distance between assets’ beta scores and the market index (S&\&&P500 for traditional assets and Bitcoin for cryptocurrencies) in each market. The smaller the distance is, the more similar the assets are to the market index.
Metrics
Stocks
(S&\&&P500 base)
US ETFs
(S&\&&P500 base)
Cryptocurrencies
(Bitcoin base)
MAE 0.272 0.156 0.178
RMSE 0.219 0.124 0.158
In summary, we found that investors in all three types of investment vehicle showed evidence of herding during the Covid-19 pandemic sub-period. This is reasonable since the serious pandemic directly affected the global economy and all financial markets. However, other sub-periods reveal different herding patterns between the cryptocurrency market and traditional markets. In particular, herding behavior was found only in stocks and US ETFs throughout 2019, while this phenomenon disappeared in the cryptocurrency market. Indeed, herding during this time appeared to be caused by the trade war between China and the US which was relevant to macroeconomic factors and directly influenced various companies in different sectors. Thus, traditional markets were undoubtedly affected the most. On the other hand, only the cryptocurrency market displayed a herding signal during the Bull market time. This can be explained by the fact that cryptocurrency investors are often swayed more by positive news and positive market returns than negative ones [88]. Whereas, traditional markets are typically more responsive to distressed market conditions than booms [117, 19]. In particular, when traditional markets experience an upward trend, professional investors are found to account for the majority portion [4]. Thus, they tend to make investment decisions by themselves which prevents them from performing as a herd. For the cryptocurrency market, however, its investors are more likely to be naïve and make irrational decisions, based on word of mouth as well as existing news on social media [57, 8]. Thus, positive news surrounding cryptocurrencies can lure inexperienced investors with the promise of substantial profits.
4.2 Herding Detection on a Combination of Cryptocurrencies, Stocks and US ETFs
In reality, investors tend to diversify their portfolios by investing in different types of assets [29]. Taking inspiration from this practice, this section combines all three types of investment vehicle used in this study to gain a more complete and realistic picture of herding behavior in financial markets, which aims to answer our second research question. However, it is unwise to conclude anything about the herding behavior of this large dataset whose assets belong to three different types of investment vehicle. Besides, there is no established approach to partition assets into distinct subsets for herding detection currently. To address this, we come up with the idea of using community detection to split the assets into smaller communities depending on their similarities. Subsequently, we apply herding detection techniques to each community. This practice is grounded in the belief that assets with common price movements are more likely to show herding behavior. Conversely, assets with different price movements tend to be used as safe havens against each other rather than being herded [118]. Thus, splitting the assets in this way can preserve the characteristics of portfolio diversification and enhance the clarity as well as granularity of herding signals while mitigating the chance of losing possible herding signals.
Furthermore, we provide information about sector distribution for each community. Specifically, each asset in a community is categorized into either cryptocurrency, or US ETF, or one of the 11 stock sectors discussed in section 3. Subsequently, the proportion of each sector is calculated in each community. Finally, for each community in each sub-period, we can find whether herding behavior exists in those assets and its corresponding sector distribution.
4.2.1 Community Structures for a Combination of Cryptocurrencies, Stocks and US ETFs
We first introduce the community structure obtained from a combination of cryptocurrencies, stocks and US ETFs in each sub-period, which is displayed in Figs 4-5. Each node in the graph represents an asset, while an edge between two nodes represents their correlation. If there is no edge between two nodes, then their correlation is negligible compared to other surrounding linked nodes. Each color represents a community in which assets with the same color belong to one community, meaning that they behave similarly and their intercorrelations are stronger than those of correlations with other nodes in different communities. The longer the path between two nodes, the more dissimilar they are.
(a) Pre-Covid-19 (04/2019 - 12/2019)
(b) Covid-19 pandemic (01/2020 - 06/2020)
(c) Post-Covid-19 (07/2020 - 12/2020)
Figure 4: Time-varying community structures of cryptocurrencies, stocks and US ETFs during: a) Pre-Covid-19, b) Covid-19 pandemic, c) Post-Covid-19. More information in each sub-figure can be seen at Clarification 3..
(a) Bull market time (01/2021 - 04/2022)
(b) Ukraine-Russia Conflict (05/2022 - 05/2023)
Figure 5: Time-varying community structures of cryptocurrencies, stocks and US ETFs during: a) Bull market period, b) Ukraine-Russia conflict. More information in each sub-figure can be seen at Clarification 3..
As can be seen in Figs 4-5, the community structures change over time, indicating the variation in correlations among assets over time. Notably, we found some common patterns out of these five different structures. Firstly, cryptocurrencies form several separate communities but these are next to each other, meaning that there is a strong correlation between cryptocurrencies themselves. Furthermore, there is still a distinction between the cryptocurrency market and traditional markets since most cryptocurrencies do not form communities with stocks and US ETFs, an example of this characteristic is highlighted within a purple-shaped oval in each figure. Secondly, it can be seen that stocks and US ETFs tend to belong to the same community. Interestingly, US ETFs seem to be the centre of a community, linked by stock assets, indicating that they are the most influential nodes in a community, an example of this characteristic is highlighted within a black-shaped oval in each figure. Thirdly, there are some stock sectors in which stocks within the same sector and ETFs that contain these stocks tend to be in the same communities, namely Energy, Healthcare, Information Technology and Finance. On the other hand, this feature does not appear in the remaining sectors. Fourthly, we also notice that S&P-related ETFs tend to belong to one community throughout the considered sub-periods.
4.2.2 Herding Behavior at Community Level
We note that due to a large number of herding results, it is challenging to compare general CSAD (Eq 3), CSAD for up and CSAD for down market conditions (Eq 4) in each community, between different communities and between different sub-periods. Therefore, we do not differentiate the results of general herding, herding during the up and herding during the down market condition, i.e. we do not go into detail about CSAD, CSAD for up and CSAD for down markets in each community. This will be explored in future works. Instead, we focus solely on determining whether the communities reveal herding behavior or not. To this end, we consider a community to exhibit herding when either CSAD or CSAD for up market condition or CSAD for down market condition reveals herding.
Figs 6-10 illustrate the herding behavior results and sector distribution of each community in each sub-period. The assets in each community are listed in Appendix E. At first glance, it is clear that the first two sub-periods look rather similar in terms of the density of herding signals, as most communities reveal herding behavior in Figs 6-7. This is contrary to the last three sub-periods (Figs 8-10) since these predominantly display a dearth of herding. Perhaps, the high density of herding during the pre-and the peak of the pandemic was caused by two major events, as mentioned previously, the US-China trade war throughout 2019 and the contagious disease Covid-19 which started early in 2020, respectively. Whereas, once the pandemic started to be controlled and the global economy recovered from the loss, financial markets resumed their normal behavior. Thus, investors stopped mimicking each other and began to make their own objective decisions, and the herding was lessened accordingly.
(a) Herding Detection Results
(b) Sector Distribution
Figure 6: Herding detection results (a) and sector distribution (b) of each community during the Pre-Covid-19 period. More information on each sub-figure can be seen at Clarification 2..
(a) Herding Detection Results
(b) Sector Distribution
Figure 7: Herding detection results (a) and sector distribution (b) of each community during the Covid-19 pandemic period. More information on each sub-figure can be seen at Clarification 2..
(a) Herding Detection Results
(b) Sector Distribution
Figure 8: Herding detection results (a) and sector distribution (b) of each community during the Post-Covid-19 period. More information on each sub-figure can be seen at Clarification 2..
(a) Herding Detection Results
(b) Sector Distribution
Figure 9: Herding detection results (a) and sector distribution (b) of each community during the Bull market period. More information on each sub-figure can be seen at Clarification 2..
(a) Herding Detection Results
(b) Sector Distribution
Figure 10: Herding detection results (a) and sector distribution (b) of each community during the Ukraine-Russia conflict period. More information on each sub-figure can be seen at Clarification 2..
We notice from Figs 6-10 that cryptocurrencies are still mostly isolated from traditional markets to some extent since they are not distributed into multiple communities. Instead, they are close to each other and tend to form 2 to 3 large communities among themselves. Observing these communities throughout different sub-periods, it is clear that they are herded most of the time. Compared to the herding detection results in the cryptocurrency market as discussed in the previous section, it is clear that without critical events impacting the entire cryptocurrency market, herding still exists in most cryptocurrencies, but at a community level, i.e herding behaviour is pronounced in different subsets of cryptocurrencies, while there was an absence of herding signals when looking at all cryptocurrencies at once, as shown in the previous section. It is challenging to find the reasons causing the herding in such cases since there was no major event that could influence the market [7, 89, 119]. However, this market seems to be mainly driven by investors’ emotions [94] and accessible news on online platforms [21]. In addition, cryptocurrency investors tend to adopt a short-term investment strategy and depend on price movements of the coins rather than reliable factors that are seen in traditional markets [90]. Consequently, the herding in those cryptocurrency communities could arise from either the potential profits or the fear of losing money at that time, encouraging investors to buy or sell. Moreover, investors might be aware of the co-movements between different cryptocurrencies. Thus, when the opportunity to gain profits (or the selling time) arrives, they tend to buy (or sell) the coins that have the same price movements to get higher profits (or mitigate the loss). This could explain why cryptocurrencies formed 2 separate communities in the Pre-Covid-19 and also the Bull market time, with both communities revealing herding behavior, while there was no herding when we detected the whole cryptocurrency market in the previous section.
Regarding other communities, each of them has a mixture of US ETFs and stocks where the US ETFs tend to act as a basket of most stocks in that community. The herding in these communities follows the same rule as the cryptocurrency communities since each sub-period reveals significant herding behavior in several communities. Remarkably, we obtained several valuable findings from the herding results of these stocks and US ETFs. In particular, there are 4 stock sectors (Energy, Healthcare, IT and Finance) in which for each sector, its assets and relevant ETFs tend to form communities among themselves, throughout the examined period. Moreover, the herding signals found in each of these sectors coincide with the timing of economic and political events that specifically influence that sector.
  • Energy Sector: In our case, the energy stocks we study mainly consist of petroleum treatment companies and petroleum exploiting companies. Thus, the herding found in this study is primarily driven by oil-related events. We found that all energy stocks and ETFs are exposed to herding across all sub-periods. Energy companies, in general, have consistently experienced tough times over the past 4 years from 2019 until the present. During this time, they were hit by a series of different events, causing large fluctuations in their corresponding stock prices over time, including the US-China trade war, the Covid-19 pandemic, the petroleum oversupply due to the shale revolution [120, 121], the Russia-Saudi Arabia price war [122, 123] and the Ukraine-Russia war [65].
  • Technology Sector: Technology stocks are grouped into several communities and reveal herding behavior in three sub-periods, including Pre-Covid-19 (coinciding with the US-China trade war [124, 125, 126, 126]), Covid-19 pandemic (coinciding with the downturn time caused by the pandemic followed by a global economic recession) and the Ukraine-Russia conflict (coinciding with the time when Technology companies were laying off workers [127, 128, 129] followed by the AI boom that gave an impetus to the Technology sector [130, 131]).
  • Healthcare Sector: Herding behavior was found in most healthcare stocks and ETFs during three sub-periods: the Covid-19 pandemic, the Post-Covid-19 and the Bull market. This herding behavior likely stems from the great attraction of investors toward pharmaceutical companies due to their positive performance during the pandemic. In particular, dozens of pharmaceutical companies began vaccine research from early 2020 due to the high demand for vaccinations for the entire world [132, 133, 134]. Moreover, other pharmaceutical companies, though not participating in producing vaccines, produced critical elements that were necessary for vaccine production such as nanoparticles or medical devices for hospitals and daily usage such as the Covid test kits. These contributions also pushed their performance during this time [134, 135].
  • Financial Sector: In this study, the financial sector mainly consists of banking firms. We found that this industry was exposed to herding behavior in the first four sub-periods, except for the Ukraine-Russia conflict time. The financial sector is strongly related to economic conditions and reflects the performance of the economy, especially the banking firms since they have the responsibility of preserving the economic stability of a country and also keeping the country away from financial crises through various schemes like reducing interest rates [136], imposing new monetary policy comes from FED [137] and cutting companies’ liquidity risks [138]. Consequently, economic uncertainty forces the financial sector to experience distress. Indeed, the global economy had been experiencing ups and downs continually through out the herding periods in this stock sector.
For the remaining sectors (i.e. Industry, Consumer Discretionary and Consumer Staples), we found that assets within the same sector never group into a large community in all sub-periods, even during the Covid-19 pandemic. Instead, they are broken down and distributed almost equally into several communities. This result indicates that investors did not perform herding on a sector basis in these sectors, even during major events. Remarkably, however, we do notice that there is simultaneous herding behavior between the Technology sector and Standard &\&& Poor (S&\&&P)-related ETFs, as shown in Fig 6 (Communities 6 and 7 for Technology stocks, Community 10 for S&\&&P-related ETFs), Fig 7 (Community 5 for Technology stocks, Community 12 for S&\&&P-related ETFs), and Fig 10 (Community 5 for Technology stocks, Community 8 for S&\&&P-related ETFs). Specifically, when the herding behavior is pronounced in the Technology sector, a set of ETFs tracking the S&P-related stock indices also shows this phenomenon, including IJR, IJS, IJT, IWM, SLYS, SLYV and SLY. These are ETFs representing the aggregate performance of small-cap stocks published by Standard & Poor. This finding is surprising since all the technology stocks used in this work are large-cap ones but their behavior is unexpectedly associated with ETFs tracking the performance of small-cap stocks. This relationship between the technology sector and these ETFs will be considered in future work.
Furthermore, we acknowledge that some communities are mixed by multiple sectors and reveal herding, but it is unclear why the herding behavior exists in those cases as there seems not to be a rule/convention for these cases. However, most of such communities appear during the economic uncertainty times, i.e. during the trade war (Fig 6) and the Covid-19 pandemic (Fig 7). By contrast, this phenomenon was not pronounced from the Bull market time onwards (Figs 9-10). Thus, it seems to be more likely that negative market conditions resulted in this phenomenon. However, since we don’t have a confident justification for this phenomenon, we will leave this result for future work.
From this experimental result, we can conclude that herding behavior should be observed from two perspectives, one from the entire market and the other from separate subsets (communities) within the market. This is because different perspectives provide us with different aspects of the market. Specifically, herding behavior occurs widely in most assets of a particular type of investment vehicle when there is a critical event that directly impacts the market. Hence, most assets involved are influenced and thus investors who are trading those assets tend to perform herding as a reaction to the event. This is why we found the herding in traditional markets during the US-China trade war, the herding in all markets during the peak of the Covid-19 pandemic and the herding in the cryptocurrency market during the Bull market time, while the remaining sub-periods show non-herding results. By contrast, we found that herding behavior occurs constantly in all types of investment vehicle but at a smaller scale, i.e. community level, as shown in Figs 6-10. In addition, herding behavior found in these communities mainly comes from specific events that solely impact that community. Subsequently, although herding might not exist in a whole market, it can be found by looking at a particular community/sector.
5 Conclusions
This work aims to detect and analyze investor herding behavior in different types of investment vehicle separately and also in their combination, including stocks, US ETFs and cryptocurrencies. We enrich the existing literature with three contributions: Firstly, the use of a more up-to-date and granular historical price time series covering several recent events, i.e. the US-China trade war in 2019, the Covid-19 pandemic followed by the global economic crisis in 2020, the Bull market period in 2021-2022 and the Ukraine-Russia conflict which started in 2022. Secondly, the use of multiple investment vehicle types adds to the absence of diversified market issues in the existing studies. Thirdly, the focus on herding behavior at a community level (small scale) enhances the clarity and granularity of this herding phenomenon.
Regarding the first research question, when examining the herding behavior in each type of investment vehicle separately, we find a similar herding pattern between stocks and US ETFs, with evidence of herding during the Pre-Covid-19 coinciding with the US-China trade war and during the Covid-19 pandemic. By contrast, the cryptocurrency market displays quite distinct behavior since its investors are seen to have performed herding during the Covid-19 pandemic, the Post-Covid-19 when the market was recovering and also during the Bull market period when assets experienced a surge in prices. In addition, the magnitude of herding in the US ETFs is much stronger than that in stocks and cryptocurrencies.
Regarding the second research question, we execute a deeper herding detection and analysis by breaking down the whole combination of stocks, cryptocurrencies and US ETFs) into smaller communities where each of these comprises the assets that are most similar to each other. Additionally, for each community, we also use information about sector distributions, with each asset belonging to one of the 13 sectors, including cryptocurrency, US ETF and 11 stock sectors. Our main finding is that herding behavior exists at all times but at a smaller scale (i.e. community level), in all types of investment vehicle examined. This mainly comes from specific events which particularly impact on that community. As a result, although herding might not be seen in a whole market composed of all assets, it can be found by looking at a particular community/subset of assets. Moreover, we notice 4 stock sectors in which for each sector, its assets and relevant ETFs tend to form communities among themselves, with the herding behavior found in these communities driven by unique events only influencing that sector. These include Energy, Healthcare, Information Technology and Finance. On the other hand, this characteristic does not appear in the remaining sectors, i.e. Industry, Consumer Discretionary and Consumer Staples.
This work enriches the understanding of herding tendencies in the examined types of investment vehicle that help investors construct an appropriate investment strategy. Furthermore, we also raise investors’ attention towards 4 stock sectors, i.e. Energy, Healthcare, Information Technology and Finance as these seem more sensitive to herding behavior than other sectors such as Industry, Consumer Discretionary and Consumer Staples. Therefore, investors who invest in these former four stock sectors should be more cautious with relevant economic and political turmoil since they are more likely to form a herding behavior during such occasions. Future work could expand on this study by using a larger dataset, covering more types of investment vehicle, such as commodities, currencies and bonds. Another area of future research could involve different time granularities. Furthermore, several findings of this study warrant further investigation: Firstly, the differences in herding magnitude between different communities. Secondly, the comparison between CSAD, CSAD for up and CSAD for down markets within each community, between different communities and between different sub-periods. Thirdly, the connections between ETFs tracking the S&\&&P-related indices and technology stocks.
Acknowledgments
The author An Pham Ngoc Nguyen wishes to acknowledge the financial support of the Science Foundation Ireland Centre for Research Training in Artificial Intelligence under grant number 18/CRT/6223. The authors Martin Crane and Marija Bezbradica wish to acknowledge the support from the Science Foundation Ireland under Grant Agreement No. 13/RC/2106_P2 at the ADAPT SFI Research Centre at DCU. ADAPT, the SFI Research Centre for AI-Driven Digital Content Technology, is funded by the Science Foundation Ireland through the SFI Research Centres Programme.
Appendix
Appendix A Figure Clarification
Clarification 1.
The process of community detection. Firstly, a set of time series is determined, where each time series corresponds to an asset. Secondly, a Pearson-based correlation matrix is constructed from the historical time series. the darker the color, the more similar the two corresponding assets. Thirdly, a distance matrix is obtained from the correlation matrix. Fourthly, a fully connected graph is constructed based on the distance between each pair of assets from the previous step. Specifically, although every two arbitrary nodes are connected, their corresponding edge has a different weight which represents the distance between them, the lower the weight, the nearer the two nodes. Fifthly, a minimum spanning tree (MST) is extracted from the distance matrix to reduce the size of the graph while keeping essential information. Lastly, Louvain community detection technique is applied to the MST to detect existing communities in the graph. Each community contains nodes that share similar characteristics and nodes in different communities have different characteristics.
Clarification 2.
Visualization of herding detection results and sector distribution. The figure illustrates herding detection results and sector distribution of each community. a) herding results for each community. The grey bar means the herding based on CSAD model (Eq 3) is significant, the green bar means the herding based on CSAD for the up market condition (Eq 4) is significant, the red bar means the herding based on CSAD for down market condition (Eq 4) is significant. The higher the bar, the more pronounced the herding; b) sector distribution (percentage of each sector) of each community, sectors are listed on the y-axis, each represented by a unique color. For each community, the longer the bar of a corresponding sector, the greater the sector’s percentage within that community.
Clarification 3.
Time-varying community structures of cryptocurrencies, stocks and US ETFs. In each sub-figure, assets with the same color belong to one community. The black-shaped oval highlights an example of a community containing stocks and US ETFs where the centre of the community is a US ETF, which is linked by surrounding stocks. Moreover, this US ETF tends to be the basket of its surrounding stocks, meaning that it is designed to track the overall performance of those stocks. The purple-shaped oval highlights communities containing cryptocurrencies. As can be seen, most cryptocurrencies are close to each other, forming communities among themselves and showing a certain distinction level from stocks and US ETFs.
Appendix B Investment Vehicles Used in this Study and Their Corresponding Assets
Table 6 introduces three types of investment vehicle used in our study and displays available assets in each type of investment vehicle.
Table 6: Three types of investment vehicle used in this study and their corresponding assets
Type Assets
Cryptocurrencies ADA (Cardano), BAT (Basic Attention Token), BCH (Bitcoin Cash), BNT (Bancor), BSV (Bitcoin SV), BTC (Bitcoin), DAI (Dai), DASH (Dash), DOGE (Dogecoin), EOS (EOS), ETH (Ethereum), ETC (Ethereum Classic), LTC (Litecoin), NEO (Neo), OMG (OMG Network), QTUM (Qtum), REP (Augur), TRX (TRON), USDT (Tether), UST (TerraUSD), XLM (Stellar), XMR (Monero), XRP (Ripple), XTZ (Tezos), XVG (Verge), ZEC (Zcash), ZRX (Ox)
US ETFs CIBR (First Trust Nasdaq Cybersecurity ETF), DIA (SPDR Dow Jones Industrial Average ETF), DJD (Invesco Dow Jones Industrial Average Dividend ETF), DJP (iPath Bloomberg Commodity Index Total Return ETN), DVY (iShares Select Dividend ETF), FDLO (Fidelity Low Volatility Factor ETF), FDN (First Trust Dow Jones Internet Index Fund), FNGU (MicroSectors FANG+ Index 3X Leveraged ETN), FTXO (First Trust Nasdaq Bank ETF), IBB (iShares Nasdaq Biotechnology ETF), IDU (iShares U.S. Utilities ETF), IEO (iShares U.S. Oil & Gas Exploration & Production ETF), IJH (iShares Core S&P Mid-Cap ETF), IJJ (iShares S&P Mid-Cap 400 Value ETF), IJK (iShares S&P Mid-Cap 400 Growth ETF), IJR (iShares Core S&P Small-Cap ETF), IJS (iShares S&P Small-Cap 600 Value ETF), IJT (iShares S&P Small-Cap 600 Growth ETF), ITB (iShares U.S. Home Construction ETF), ITOT (iShares Core S&P Total U.S. Stock Market ETF), IUSG (iShares Core S&P U.S. Growth ETF), IVE (iShares S&P 500 Value ETF), IWB (iShares Russell 1000 ETF), IWM (iShares Russell 2000 ETF), IWV (iShares Russell 3000 ETF), IYE (iShares U.S. Energy ETF), IYF (iShares U.S. Financials ETF), IYJ (iShares U.S. Industrials ETF), IYR (iShares U.S. Real Estate ETF), IYT (iShares Transportation Average ETF), IYW (iShares U.S. Technology ETF), IYY (iShares Dow Jones U.S. ETF), KBWB (Invesco KBW Bank ETF), OEF (iShares S&P 100 ETF), ONEQ (Fidelity Nasdaq Composite ETF), PNQI (Invesco Nasdaq Internet ETF), QQQ (Invesco QQQ Trust), QTEC (First Trust NASDAQ-100 Technology Sector Index Fund), QYLD (Global X NASDAQ 100 Covered Call ETF), ROBT (First Trust Nasdaq Artificial Intelligence and Robotics ETF), RPG (Invesco S&P 500 Pure Growth ETF), RPV (Invesco S&P 500 Pure Value ETF), SLYG (SPDR S&P 600 Small Cap Growth ETF), SLYV (SPDR S&P 600 Small Cap Value ETF), SLY (SPDR S&P 600 Small Cap ETF), SOXX (iShares PHLX Semiconductor ETF), SPY (SPDR S&P 500 ETF), SVXY (ProShares Short VIX Short-Term Futures ETF), XTN (SPDR S&P Transportation ETF).
Stocks AAPL (Apple Inc.), ABBV (AbbVie Inc.), ABB (ABB Ltd.), ABT (Abbott Laboratories), ACN (Accenture plc), ADBE (Adobe Inc.), ADI (Analog Devices, Inc.), ADP (Automatic Data Processing, Inc.), AIR (AAR Corp.), ALV (Autoliv Inc.), AMAT (Applied Materials, Inc.), AMD (Advanced Micro Devices, Inc.), AMGN (Amgen Inc.), AMT (American Tower Corporation), AMZN (Amazon.com Inc.), ASML (ASML Holding NV), ASX (ASE Technology Holding Co., Ltd.), AVGO (Broadcom Inc.), AXP (American Express Company), AZN (AstraZeneca PLC), BABA (Alibaba Group Holding Limited), BAC (Bank of America Corp.), BA (The Boeing Company), BDX (Becton, Dickinson and Company), BHP (BHP Group Limited), BKNG (Booking Holdings Inc.), BKR (Baker Hughes Company), BLK (BlackRock Inc.), BMY (Bristol Myers Squibb Company), BP (BP plc), BSX (Boston Scientific Corporation), BTI (British American Tobacco plc), BUD (Anheuser-Busch InBev SA/NV), BX (The Blackstone Group Inc.), CAT (Caterpillar Inc.), CB (Chubb Limited), CFR (Cullen/Frost Bankers, Inc.), CI (Cigna Corporation), CMCSA (Comcast Corporation), CNI (Canadian National Railway Company), COP (ConocoPhillips), COST (Costco Wholesale Corporation), CP (Canadian Pacific Railway Limited), CRM (Salesforce.com Inc.), CSCO (Cisco Systems, Inc.), CSL (Carlisle Companies Incorporated), CVS (CVS Health Corporation), CVX (Chevron Corporation), C (Citigroup Inc.), DEO (Diageo plc), DE (Deere & Company), DHR (Danaher Corporation), DIS (The Walt Disney Company), DTE (DTE Energy Company), ELV (Eaton Vance Corporation), EL (The Estée Lauder Companies Inc.), EOG (EOG Resources Inc.), EQNR (Equinor ASA), ETN (Eaton Corporation plc), GE (General Electric Company), GILD (Gilead Sciences, Inc.), GOOG (Alphabet Inc.), GS (The Goldman Sachs Group, Inc.), HCA (HCA Healthcare, Inc.), HDB (HDFC Bank Limited), HD (The Home Depot, Inc.), HON (Honeywell International Inc.), HSBC (HSBC Holdings plc), IBM (International Business Machines Corporation), IBN (ICICI Bank Limited), IDEX (Ideanomics, Inc.), INTC (Intel Corporation), INTU (Intuit Inc.), ISRG (Intuitive Surgical, Inc.), ITW (Illinois Tool Works Inc.), JNJ (Johnson & Johnson), JPM (JPMorgan Chase & Co.), KO (The Coca-Cola Company), LIN (Linde plc), LLY (Eli Lilly and Company), LMT (Lockheed Martin Corporation), LOW (Lowe’s Companies, Inc.), LRCX (Lam Research Corporation), MA (Mastercard Incorporated), MCD (McDonald’s Corporation), MC (Moelis & Company), MDLZ (Mondelez International, Inc.), MDT (Medtronic plc), META (Meta Platforms, Inc.), MMC (Marsh & McLennan Companies, Inc.), MO (Altria Group, Inc.), MRK (Merck & Co., Inc.), MSFT (Microsoft Corporation), MS (Morgan Stanley), MUFG (Mitsubishi UFJ Financial Group, Inc.), MU (Micron Technology, Inc.), NEE (NextEra Energy, Inc.), NFLX (Netflix, Inc.), NKE (NIKE, Inc.), NOW (ServiceNow, Inc.), NVDA (NVIDIA Corporation), NVO (Novo Nordisk A/S), NVR (NVR, Inc.), NVS (Novartis AG), ORCL (Oracle Corporation), PANW (Palo Alto Networks, Inc.), PBR (Petroleo Brasileiro S.A. - Petrobras), PDD (Pinduoduo Inc.), PEP (PepsiCo, Inc.), PFE (Pfizer Inc.), PG (Procter & Gamble Co.), PLD (Prologis, Inc.), PM (Philip Morris International Inc.), PYPL (PayPal Holdings, Inc.), QCOM (QUALCOMM Incorporated), SAP (SAP SE), SBUX (Starbucks Corporation), SCHW (The Charles Schwab Corporation), SHEL (Sealed Air Corporation), SHOP (Shopify Inc.), SLB (Schlumberger Limited), SNY (Sanofi), SONY (Sony Group Corporation), SO (The Southern Company), SYK (Stryker Corporation), TD (The Toronto-Dominion Bank), TJX (The TJX Companies, Inc.), TMO (Thermo Fisher Scientific Inc.), TMUS (T-Mobile US, Inc.), TM (Toyota Motor Corporation), TSLA (Tesla, Inc.), TSM (Taiwan Semiconductor Manufacturing Company Limited), TTE (TotalEnergies SE), TXN (Texas Instruments Incorporated), T (AT&T Inc.), UL (Unilever PLC), UNH (UnitedHealth Group Incorporated), UNP (Union Pacific Corporation), UPS (United Parcel Service, Inc.), VRTX (Vertex Pharmaceuticals Incorporated), VZ (Verizon Communications Inc.), V (Visa Inc.), WFC (Wells Fargo & Co.), WMT (Walmart Inc.), XOM (Exxon Mobil Corporation), ZTS (Zoetis Inc.)
Appendix C Classification of Financial Sectors in the Stock Market
Table 7 lists all stock sectors in the stock market and available assets from our dataset in each sector.
Table 7: Distribution of Stock Sectors in the Stock Dataset. Each Asset is Assigned to an Appropriate Stock Sector
Information
Technology
Healthcare Financials
Consumer
discretionary
Consumer
staples
Industrials Energy Utilities Materials Real estate
Communication
services
ASX
AAPL
ACN
ADBE
ADI
ADP
AMAT
AMD
ASML
AVGO
BABA
CMCSA
CRM
CSCO
GOOG
IBM
INTC
INTU
LRCX
MA
META
MSFT
MU
NFLX
NOW
NVDA
ORCL
PANW
PDD
QCOM
SAP
SHOP
SONY
TMUS
TSM
TXN
T
VZ
V
ABBV
ABT
AMGN
AZN
BDX
BMY
BSX
CI
CVS
DHR
ELV
GILD
HCA
ISRG
JNJ
LLY
MDT
MRK
NVO
NVS
PFE
SNY
SYK
TMO
UNH
VRTX
ZTS
AXP
BAC
BLK
BX
CB
CFR
C
GS
HDB
HSBC
IBN
JPM
MC
MMC
MS
MUFG
SCHW
TD
WFC
ALV
AZMN
BKNG
COST
CSL
DIS
HD
IDEX
LOW
MCD
NKE
NVR
SBUX
TJX
TM
TSLA
BTI
BUD
DEO
EL
GE
HON
ITW
KO
MDLZ
MO
PEP
PG
PM
UL
WMT
ABB
AIR
BA
CAT
CNI
CP
DE
ETN
LMT
UNP
UPS
PYPL
BKR
BP
COP
CVX
EQNR
PBR
SHEL
SLB
TTE
EOG
XOM
DTE
NEE
SO
BHP
LIN
AMT
PLD
Appendix D Details of Assets in each Community (Entire Period)
Table 8 provides the result of community detection on the combination of stocks, cryptocurrencies and US ETFs throughout the entire period from 04/2019 until 05/2023.
Table 8: Assets in each community over the entire period. The red color indicates cryptocurrencies, the blue color indicates US ETFs and the black color indicates stocks
Community Assets
1 AAPL, ADBE, AMGN, GILD, GOOG, INTU, MSFT, SONY, TSLA, VRTX, FNGU, IBB, IYW, ONEQ, QQQ, QYLD
2 ABBV, AMT, BMY, BSX, CMCSA, ETN, HCA, HON, IBM, ITW, KO, LLY, MDLZ, MDT, MMC, PEP, PG, PLD, SYK, DJD, IVE, IYJ, IYR
3 ABB, ACN, BLK, BX, DIS, EL, ISRG, MA, NKE, ORCL, SAP, SBUX, TMUS, V, ZTS, FDLO, ITOT, IWB, IWV, IYY, SPY, SVXY
4 BTI, CAT, CB, DE, DTE, LMT, MO, NEE, PM, SO, T, VZ, DVY, IDU
5 ADI, AMAT, AMD, ASML, ASX, AVGO, INTC, LRCX, MU, NVDA, QCOM, TSM, TXN, QTEC, ROBT, SOXX
6 ALV, BHP, BKNG, CNI, CP, CSL, MC, NVR, TJX, UNP, UPS, IJH, IJJ, IJK, ITB, IYT, XTN
7 ABT, BDX, COST, DHR, TMO, WMT, BNT, BTC, DOGE, ETH, USDT, XVG, IUSG, OEF, RPG
8 AIR, IJR, IJS, IJT, IWM, SLYG, SLYV, SLY
9 AMZN, BABA, CRM, IDEX, META, NFLX, NOW, PANW, PDD, PYPL, SHOP, CIBR, FDN, PNQI
10 ADA, BAT, BCH, BSV, DAI, DASH, EOS, ETC, LTC, NEO, OMG, QTUM, REP, TRX, XLM, XMR, XRP, XTZ, ZEC, ZRX
11 ADP, AZN, BA, BUD, CI, CSCO, CVS, DEO, ELV, HD, JNJ, LIN, LOW, MCD, MRK, NVO, NVS, PFE, SNY, TM, UL, UNH, DIA
12 AXP, BAC, CFR, C, GS, HDB, HSBC, IBN, JPM, MS, MUFG, SCHW, TD, WFC, FTXO, IYF, KBWB
13 BKR, BP, COP, CVX, EOG, EQNR, GE, PBR, SHEL, SLB, TTE, XOM, DJP, IEO, IYE, RPV
Appendix E Details of Assets in each Community (Sub-Period)
Tables 9 - 13 provide the results of community detection on the combination of stocks, cryptocurrencies and US ETFs during different sub-periods, including Pre-Covid-19 (04/2019 - 12/2019), Covid-19 pandemic (01/2020 - 06/2020), Post-Covid-19 (07/2020 - 12/2020), Bull market time (01/2021 - 04/2022), Ukraine-Russia conflict (05/2022 - 05/2023).
Table 9: Assets in each community during the Pre-Covid-19. The red color indicates cryptocurrencies, the blue color indicates US ETFs and the black color indicates stocks
Community Assets
1 ACN, ADP, AMT, DTE, EL, INTU, ISRG, LMT, MA, MCD, NEE, NOW, PLD, PYPL, SAP, SBUX, SONY, SO, V, IDU, IUSG, IYR
2 ABBV, AXP, BUD, CMCSA, COST, HON, LIN, MMC, TJX, WMT, OEF, SPY, SVXY
3 ABB, BHP, BLK, CAT, CB, CI, CVS, DE, ELV, ETN, GE, HSBC, ITW, MUFG, TD, UNH, IVE, IYF
4 ABT, BDX, BSX, DHR, HDB, IBN, LLY, MDT, MRK, NVS, PFE, SYK, TMO, ZTS
5 BAT, BNT, DAI, DASH, ETH, OMG, USDT, UST, XLM, XMR, XRP, XVG, ZEC, ZRX
6 ADBE, CRM, GOOG, META, PANW, SHOP, CIBR, FDN
7 ADI, AMAT, AMD, ASX, AVGO, INTC, LRCX, MU, NVDA, QCOM, TSM, TXN, SOXX
8 AIR, BX, CNI, CP, CSL, MC, NVR, ORCL, TMUS, UNP, UPS, IJH, IJJ, IJK, ITB, ITOT, IYT, XTN
9 ALV, BAC, CFR, C, GS, JPM, MS, SCHW, WFC, FTXO, KBWB, RPV
10 AMGN, BMY, GILD, HCA, VRTX, IBB, IJR, IJS, IJT, IWM, SLYG, SLYV, SLY
11 AZN, DEO, KO, MDLZ, NVO, PEP, PG, SNY, TM, UL, FDLO, IWB, IWV, IYJ, IYY, RPG
12 ADA, BCH, BSV, BTC, DOGE, EOS, ETC, LTC, REP
13 AAPL, AMZN, ASML, BABA, BKNG, CSCO, IDEX, MSFT, NFLX, PDD, TSLA, FNGU, IYW, ONEQ, PNQI, QQQ, QTEC, QYLD, ROBT
14 BA, DIS, HD, IBM, JNJ, LOW, NKE, DIA, DJD
15 BKR, BP, COP, CVX, EOG, EQNR, PBR, SHEL, SLB, TTE, XOM, DJP, IEO, IYE
16 BTI, MO, PM, T, VZ, NEO, QTUM, TRX, XTZ, DVY
Table 10: Assets in each community during the Covid-19 pandemic. The red color indicates cryptocurrencies, the blue color indicates US ETFs and the black color indicates stocks
Community Assets
1 AAPL, ADBE, AMGN, AMZN, COST, GILD, GOOG, INTC, MSFT, PYPL, SONY, TMUS, VRTX, WMT, IBB, QQQ, QYLD
2 BTI, CB, DE, DTE, ETN, MO, NEE, SO, T, VZ, DVY, IDU
3 ABB, BUD, CI, DEO, DIS, ELV, EL, HD, LOW, MCD, NKE, SBUX, UNH, DIA
4 ABBV, ABT, AZN, BDX, CSCO, DHR, IDEX, JNJ, MA, MRK, NVO, NVS, ORCL, PFE, SNY, TMO, V, IUSG, OEF, ONEQ, SVXY
5 ADI, AMAT, AMD, ASML, ASX, AVGO, LRCX, MU, NVDA, QCOM, TSM, TXN, SOXX
6 ADP, BSX, CAT, CMCSA, CNI, CP, CVS, GE, HON, IBM, ITW, KO, LMT, PM, UNP, UPS, DJD, IVE, IYJ, IYT, XTN
7 INTU, NOW, ADA, BAT, BNT, ETH, NEO, OMG, QTUM, XMR, XTZ, XVG, ZRX, IYW, QTEC, RPG
8 ALV, BA, BHP, BKR, BP, COP, CVX, EOG, EQNR, MUFG, PBR, SHEL, SLB, TD, TTE, XOM, DJP, IEO, IYE, RPV
9 AMT, AXP, BAC, BX, CFR, C, GS, HDB, HSBC, IBN, JPM, MS, PLD, SCHW, TJX, WFC, FTXO, IYF, IYR, KBWB
10 BABA, CRM, META, NFLX, PANW, PDD, SHOP, TSLA, CIBR, FDN, FNGU, PNQI
11 BCH, BSV, BTC, DAI, DASH, DOGE, EOS, ETC, LTC, REP, TRX, XLM, XRP, ZEC
12 AIR, BKNG, CSL, HCA, MC, NVR, TM, IJH, IJJ, IJK, IJR, IJS, IJT, ITB, IWM, ROBT, SLYG, SLYV, SLY
13 ACN, BLK, BMY, ISRG, LIN, LLY, MDLZ, MDT, MMC, PEP, PG, SAP, SYK, UL, ZTS, USDT, UST, FDLO, ITOT, IWB, IWV, IYY, SPY
Table 11: Assets in each community during the Post-Covid-19. The red color indicates cryptocurrencies, the blue color indicates US ETFs and the black color indicates stocks
Community Assets
1 BX, ISRG, FDLO, ITOT, IUSG, IWB, IWV, IYY, OEF, ROBT, RPG
2 ABBV, AZN, BMY, GILD, JNJ, LLY, MRK, NVO, NVS, PFE, SNY
3 ABB, ABT, ACN, BDX, BLK, DHR, HD, LOW, MA, ORCL, TMO, V, ZTS, SPY, SVXY
4 ADBE, CRM, META, NFLX, NOW, PANW, PDD, PYPL, SHOP, CIBR, FDN, PNQI
5 ADI, AMAT, ASML, ASX, AVGO, INTC, LRCX, MU, QCOM, TSM, TXN, SOXX
6 ADP, BSX, CAT, CMCSA, CSCO, CVS, DE, ETN, HON, IBM, ITW, KO, LMT, MDLZ, MDT, MMC, PEP, PG, SYK, TJX, DJD, IVE
7 AIR, BA, CSL, GE, HCA, MC, MUFG, TM, IJJ, IJS, RPV, SLYV
8 CI, DIS, ELV, EL, HDB, IBN, LIN, MCD, NKE, SAP, SBUX, UNH, DIA
9 ALV, CNI, CP, NVR, UNP, UPS, IJH, IJK, IJR, IJT, ITB, IWM, IYJ, IYT, SLYG, SLY, XTN
10 AMGN, AMZN, BABA, COST, GOOG, INTU, SONY, TMUS, TSLA, VRTX, WMT, FNGU, IBB, ONEQ, QQQ, QYLD
11 AMT, AXP, BKNG, GS, MS, PLD, IYF, IYR
12 BAC, CFR, C, JPM, SCHW, WFC, DAI, FTXO, KBWB
13 IDEX, BAT, BCH, BSV, DOGE, EOS, ETC, NEO, OMG, QTUM, REP, TRX, XLM, XRP, XTZ, XVG, ZRX
14 AAPL, AMD, MSFT, NVDA, ADA, BNT, BTC, DASH, ETH, LTC, XMR, ZEC, IYW, QTEC
15 BHP, BKR, BP, BUD, COP, CVX, DEO, EOG, EQNR, PBR, SHEL, SLB, TTE, UL, XOM, DJP, IEO, IYE
16 BTI, CB, DTE, HSBC, MO, NEE, PM, SO, TD, T, VZ, USDT, UST, DVY, IDU
Table 12: Assets in each community during the Bull market time. The red color indicates cryptocurrencies, the blue color indicates US ETFs and the black color indicates stocks
Community Assets
1 AAPL, ADBE, AMZN, GOOG, INTU, MSFT, TSLA, FNGU, IYW, ONEQ, QQQ, QTEC, QYLD
2 AMT, BTI, CMCSA, CVS, DTE, ETN, HCA, ITW, KO, MO, NEE, PLD, PM, SO, DAI, DJD, IDU, IVE, IYR
3 CNI, CP, MA, UNP, UPS, V, IYJ, IYT, XTN
4 CAT, DE, GE, IBM, LMT, T, VZ, DVY, IJJ, RPV
5 ADI, AMAT, AMD, ASML, ASX, AVGO, INTC, LRCX, MU, NVDA, QCOM, TSM, TXN, SOXX
6 AIR, ALV, BA, CSL, HD, IDEX, LOW, MC, NVR, IJH, IJK, IJR, IJS, IJT, ITB, IWM, SLYG, SLYV, SLY
7 ABT, BDX, DHR, ISRG, SAP, TMO, VRTX, ZTS, IBB, IUSG, OEF, ROBT, RPG
8 ABBV, AMGN, AZN, BMY, GILD, JNJ, LLY, MDLZ, MRK, NVO, NVS, PEP, PFE, PG, SNY, UL, UST
9 ABB, ACN, ADP, BLK, BSX, COST, EL, MDT, ORCL, SYK, TMUS, WMT, FDLO,SPY, SVXY
10 AXP, BAC, CB, CFR, C, GS, HSBC, JPM, MS, MUFG, SCHW, WFC, FTXO, IYF, KBWB
11 BABA, CRM, META, NFLX, NOW, PANW, PDD, PYPL, SHOP, CIBR, FDN, PNQI
12 BHP, BKR, BP, COP, CVX, EOG, EQNR, PBR, SHEL, SLB, TTE, XOM, DJP, IEO, IYE
13 BKNG, BUD, CI, CSCO, DEO, ELV, HON, LIN, MCD, MMC, SBUX, TJX, TM, UNH, DIA
14 TD, ADA, BAT, BCH, BSV, DASH, EOS, ETC, NEO, OMG, QTUM, REP, TRX, USDT, XLM, XMR, XRP, XTZ, ZEC
15 BX, DIS, HDB, IBN, NKE, SONY, ITOT, IWB, IWV, IYY
16 BNT, BTC, DOGE, ETH, LTC, XVG, ZRX
Table 13: Assets in each community during the Ukraine-Russia Conflict. The red color indicates cryptocurrencies, the blue color indicates US ETFs and the black color indicates stocks
Community Assets
1 BHP, BKR, BP, COP, CVX, EOG, EQNR, PBR, SHEL, SLB, TTE, XOM, DJP, IEO, IYE
2 ABBV, BMY, JNJ, KO, LLY, MDLZ, MO, MRK, PEP, PFE, PG, PM
3 AMT, BTI, CAT, CB, CMCSA, DE, DTE, ITW, LMT, MMC, NEE, PLD, SO, T, VZ, DVY, IDU, IVE, IYR, RPV
4 ABT, ACN, BDX, BLK, BSX, BX, DIS, EL, HDB, IBN, ISRG, MA, MDT, NKE, SAP, SBUX, SONY, SYK, V, ITOT, SPY, SVXY
5 ADBE, ADI, AMAT, AMD, ASML, ASX, AVGO, GOOG, IDEX, INTC, LRCX, MSFT, MU, NVDA, QCOM, TSM, TXN, IYW, QTEC, ROBT, SOXX
6 BUD, CI, COST, CSCO, CVS, DEO, ELV, HON, IBM, LIN, MCD, TMUS, UL, UNH, WMT, DIA, DJD
7 ADP, ORCL, TM, ZTS, FDLO, IWB, IWV, IYJ, IYY, RPG
8 AIR, IJR, IJS, IJT, IWM, SLYG, SLYV, SLY
9 AMGN, AZN, GILD, NVO, NVS, SNY, VRTX, IBB
10 ABB, ALV, BA, BKNG, CNI, CP, CSL, ETN, GE, HCA, HD, LOW, MC, NVR, TJX, UNP, UPS, IJH, IJJ, IJK, ITB, IYT, XTN
11 AMZN, CRM, NFLX, NOW, PANW, PYPL, SHOP, CIBR, FDN, PNQI
12 AXP, BAC, CFR, C, GS, HSBC, JPM, MS, MUFG, SCHW, TD, WFC, DAI, FTXO, IYF, KBWB
13 ADA, BCH, BNT, BSV, BTC, DASH, DOGE, EOS, ETC, ETH, LTC, NEO, OMG, QTUM, TRX, USDT, XLM, XMR, XRP, XVG, ZEC
14 AAPL, BABA, DHR, INTU, META, PDD, TMO, TSLA, BAT, REP, XTZ, ZRX, FNGU, IUSG, OEF, ONEQ, QQQ, QYLD